15:30
1.5 MeV/u beam currently being tuned
Target is ON, setpoint = 135 K
System checks OK - white rabbit fail diff ~ 9
Temperatures OK
Merger OK
Data rate ~ 1400 kb/sec
Leakage current ~ 3.41 uA
Aida 1 and 2 missing channels from previous days appear working on stat layout 2 and xy pixel total plot (attachment 3)
17:12
Beam tuning complete. See attached Ex rate histogram.
The cryo compressor is being turned OFF for around 1 hour to remove any frozen nitrogen which may be the cause of the increased pressures observed.
Beam OFF to do so.
System checks OK - white rabbit fail diff ~ 9
Temperatures OK
Merger OK
Data rate ~ 200 kb/sec
Compressor OFF, setpoint on controller is 273 K
18:51
Pressure in the ring recovered quickly after closing valves to gas jet (attachment 5)
Increasing temperature caused a large spike in pressure in the inlets. E1 turbo briefly shut down due to the pressure increase, since restarted and pressure back in the -4 mbar range. (attachment 6)
Temperature being decreased
Beam back ON
data rate ~ 1500 kb/sec
merger OK
Temperatures OK
System checks OK - white rabbit diff ~ 12
21:58
After cooling the jet target, expected density is not achieved. E4 pressure is ~4E-10 mbar compared to ~2E-8mbar previously. Temperature reduced to 127K E4 pressure increases however stilll within -10 mbar range.
With beam on target we see rutherford peak at ~3 MeV, however it is much reduced.
Beam OFF to investigate inside the cave.
Decreasing temperature of nozzle in 0.5 k steps to see effect on E4 and S1 dump.
|